# AdaFold: Adapting Folding Trajectories of Cloths via Feedback-loop Manipulation

### v1

@@ VII-B Folding Optimization Results
The goal of this section is two-fold: assessing the relevance of optimizing the folding trajectory and showcasing the benefits of feedback-loop optimization. We conducted the experiment in simulation, and we compared our method against the four baselines presented in the previous section. We measure the folding success using the IoU metric. We evaluated all the methods under two conditions: one where the folding task utilized the same cloth as in training and another where $10$ cloth parameters were randomly selected to introduce variations not observed during training. We repeated the folding 20 times for each combination of cloth parameters and method.
@@
Table II shows the folding results. It can be observed that methods that do not optimize the folding trajectory, specifically the Random and Fixed baselines, yielded the lowest performance, thereby underscoring the advantage of trajectory optimization. The Random baselines had, on average, a similar performance to the Triangular baseline, as our constrained sampling approach favours some alignment at the end of the execution by completing the random trajectory till the place position. While the model free-baselines (DDPG-Critic) improved the fold compared to the Triangular baseline, the small amount of offline data proved to be a challenging scenario for learning a good value function for folding. A similar result can be observed for the open-loop approach (AdaFold-OL), which improved the fold compared to the fixed trajectory but obtained, on average, a worse result than the feedback-loop variant. AdaFold outperformed all the baselines and achieved the best folds, showcasing the benefits of feedback-loop optimization. Finally, the small difference between the results over $1$ or $N$ cloths of the Triangular baseline suggests that the variation in cloth behaviours that can be simulated by PyBullet is not large, as studied in [37]. Thus, while AdaFold folding success for $N$ cloths still outperformed the Triangular trajectory, we defer the evaluation of AdaFold’s generalization to cloths with varying physical properties to the real-world experiments.
TABLE III: Final IoU evaluated on real-world cloths $2-6$. The folding is repeated 5 times for each combination of cloth and method.
Cloth Triangular $\uparrow$ AdaFold $\uparrow$
2 $0.57\pm 0.03$ $0.81\pm 0.06$
3 $0.55\pm 0.03$ $0.81\pm 0.06$
4 $0.61\pm 0.01$ $0.79\pm 0.04$
5 $0.75\pm 0.02$ $0.83\pm 0.04$
6 $0.60\pm 0.03$ $0.82\pm 0.05$
All $0.62\pm 0.08$ $\mathbf{0.81\pm 0.05}$
@@
Method 10 Random Poses $\uparrow$ Cloth 1 $\uparrow$
Triangular $0.57\pm 0.03$ $0.55\pm 0.01$
AdaFold $\mathbf{0.76\pm 0.09}$ $\mathbf{0.71\pm 0.05}$
@@ VII-D Real World Experiments
This evaluation investigates to what extent AdaFold generalizes to different cloth variations including physical properties, position and size. First, we evaluated variations in physical properties by assessing whether AdaFold can improve the fold quality compared to the Triangular trajectory for cloths $2-6$. We evaluated the folding outcome with the IoU metric. We present the results in Table III. The variance that the Triangular trajectory exhibits highlights that identical folding trajectories can yield different outcomes based on the cloth’s properties, where stiffer cloths (e.g. samples $5-6$) achieved on average a better fold with respect to less stiff cloths. Conversely, AdaFold consistently produced satisfactory folds across different samples, improving the fold for samples where the fixed trajectory fell short. Figure 1 provides a qualitative comparison of the fixed and optimized trajectories for samples $2$ and $6$.
We extended this evaluation by testing AdaFold on variations in the initial positions and shapes of the cloth. For the former variation, we selected cloth $2$ and randomly rotated its starting position on the table $10$ times, with rotations ranging between ±$\pm 45^{\circ}$ degrees. For the latter variation, we selected cloth $1$, which is larger than clothes $2-6$. We compared AdaFold against the Triangular baseline. The results, detailed in Table IV, show AdaFold outperforming the Triangular trajectory in both scenarios. The overall outcome confirms the adaptability of our method, underscoring the benefit of using particle-based representations along feedback-loop manipulation to adapt to different object variations. Nevertheless, the performance observed with these last two variations was not as high as in earlier evaluations, possibly due to the higher misalignment between training and test conditions.
Figure 6: Visualization of the effect of our filtering approach. Wrong descriptors extracted with the Greedy selection are successfully obtained with the Ensemble selection.
@@
Camera(s) No Semantic $\downarrow$ Greedy $\downarrow$ Ensemble $\downarrow$
Front $0.40\pm 0.17$ $0.21\pm 0.20$ $0.10\pm 0.11$
Back $0.29\pm 0.15$ $0.20\pm 0.21$ $0.07\pm 0.11$
Both $0.32\pm 0.14$ $0.22\pm 0.21$ $\mathbf{0.04\pm 0.08}$
@@ VII-E Semantic Cloth Representation
Finally, we assess whether our proposed perception module improves the representation of the cloth. Unlike previous sections, now the IoU represents a quantitative description of the cloth state instead of a performance metric. We used a subset of the real-world trajectories recorded for the previous evaluations as a dataset, and we manually annotated the mask belonging to the upper or bottom half of the cloth to extract $P^{U}$ and $P^{B}$ and compute the ground truth IoU. To compare different state representations, we used the mean absolute error (MAE) computed between the estimated IoU from a specific cloth representation and the ground truth. Specifically, we compared the representations with semantic descriptors obtained through our mask selection approach (Ensemble), against those derived from a greedy selection of the highest confidence mask from the VLM (Greedy), and a representation lacking semantic descriptors (No Semantic). We further integrated the comparison between one and two camera points of view. As shown in Table V, our approach always achieved the lowest MAE, with the best performance provided by leveraging the observations from both cameras. The representation with no semantic information obtained the worst results, confirming the challenges of distinguishing different folded states only relying on the observed points of the cloth. While integrating semantic information reduced the MAE, blindly trusting the VLM (Greedy) was still prone to error, as the pre-trained model failed to segment the cloth. On the other hand, our simple filtering technique improved the quality of the representation due to a more accurate segmentation, as visualized in Fig. 6. These results confirm that our proposed state representation better describes different folded configurations, and our filtering technique provides an accurate segmentation.
@@ VIII Conclusion and Future Work
We introduced AdaFold, a novel framework that leverages model-based feedback-loop manipulation to optimize cloth folding trajectories. AdaFold addresses some of the perception challenges by integrating semantic descriptors extracted from pre-trained VLMs into point cloud representations. Our experiments validated the hypothesis that AdaFold adapts folding trajectories to variations in physical properties, positions, and sizes of cloths. With our half folding task we showcased the potential of coupling a strong perception module with data-driven optimization strategies to perform feedback-loop manipulation. We plan to extend these results to different goal and reward configurations, a broader spectrum of clothing items, and tasks beyond folding.

### v3

@@ VII-B Folding Optimization Results
The goal of this section is two-fold: assessing the relevance of optimizing the folding trajectory and showcasing the benefits of feedback-loop optimization. We conducted the experiment in simulation, and we compared our method against the five baselines presented in the previous section. We measured the folding success using the IoU metric. We evaluated all the methods under two conditions: one where the folding task utilized the same cloth as in training and another where $10$ cloth parameters were randomly selected to introduce variations not observed during training. We repeated the folding 20 times for each combination of cloth parameters and method.
@@
Table II shows the folding results. It can be observed that methods that do not optimize the folding trajectory, specifically the Random and Fixed baselines, yielded the lowest performance, thereby underscoring the advantage of trajectory optimization. The Random baselines had, on average, a similar performance to the Triangular baseline, as our constrained sampling approach favours some alignment at the end of the execution by completing the random trajectory till the place position. While the model free-baselines (DDPG-Critic) improved the fold compared to the Triangular baseline, the small amount of offline data proved to be a challenging scenario for learning a good value function for folding. A similar result can be observed for AdaFold’s ablations (AdaFold-OL, AdaFold-NL), which improved the fold compared to the fixed trajectory but obtained, on average, a worse result than the feedback-loop variant. AdaFold outperformed all the baselines and achieved the best folds, showcasing the benefits of feedback-loop optimization. Finally, the small difference between the results over $1$ or $N$ cloths of the Triangular baseline suggests that the variation in cloth behaviours that can be simulated by PyBullet is not large, as studied in [38]. Thus, while AdaFold folding success for $N$ cloths still outperformed the Triangular trajectory, we defer the evaluation of AdaFold’s generalization to cloths with varying physical properties to the real-world experiments.
TABLE III: Final IoU evaluated on real-world Cloths (C.) $2-6$. The folding is repeated 5 times for each combination of cloth and method. Best viewed with zoom.
[…]
2 $0.63\pm 0.03$ $0.39\pm 0.10$ $0.65\pm 0.18$ $0.51\pm 0.07$ $\mathbf{0.72\pm 0.07}$
3 $0.53\pm 0.04$ $0.33\pm 0.06$ $0.61\pm 0.08$ $0.58\pm 0.04$ $\mathbf{0.65\pm 0.06}$
4 $0.61\pm 0.01$ $0.35\pm 0.18$ $0.70\pm 0.06$ $0.63\pm 0.02$ $\mathbf{0.74\pm 0.04}$
5 $0.72\pm 0.09$ $0.37\pm 0.02$ $0.52\pm 0.12$ $0.59\pm 0.03$ $\mathbf{0.74\pm 0.06}$
6 $0.75\pm 0.11$ $0.45\pm 0.16$ $0.77\pm 0.09$ $0.64\pm 0.03$ $\mathbf{0.81\pm 0.05}$
@@
Method 10 Poses $\uparrow$ Cloth 1 $\uparrow$ Leg $\uparrow$ Pants $\uparrow$
Triangular $0.63\pm 0.03$ $0.64\pm 0.03$ $0.58\pm 0.05$ $0.64\pm 0.03$
AdaFold $\mathbf{0.72\pm 0.08}$ $\mathbf{0.76\pm 0.04}$ $\mathbf{0.64\pm 0.04}$ $\mathbf{0.68\pm 0.05}$
@@ VII-D Real World Experiments
This evaluation investigates to what extent AdaFold generalizes to different cloth variations including physical properties, position, size and shape.
Physical properties: First, we evaluated variations in physical properties by assessing whether AdaFold can improve the fold quality compared to the Triangular trajectory for cloths $2-6$. We evaluated the folding outcome with the IoU metric. We present the results in Table III. The variance that the Triangular trajectory exhibits highlights that identical folding trajectories can yield different outcomes based on the cloth’s properties, where stiffer cloths (e.g. samples $5-6$) achieved on average a better fold with respect to less stiff cloths. Conversely, AdaFold consistently produced satisfactory folds across different samples, improving the fold for samples where the fixed trajectory fell short. Figure 1 provides a qualitative comparison of the fixed and optimized trajectories for samples $2$ and $6$.
Position and size: We extended this evaluation by testing AdaFold on variations in the initial positions and shapes of the cloth. For the former variation, we selected cloth $2$ and randomly rotated its starting position on the table $10$ times, with rotations ranging between ±$\pm 45^{\circ}$ degrees. For the latter variation, we selected cloth $1$, which is larger than clothes $2-6$. We compared AdaFold against the Triangular baseline. The results, detailed in Table IV under the columns “10 Poses” and “Cloth 1”, show AdaFold outperforming the Triangular trajectory in both scenarios.
Shape: We finally tested AdaFold’s ability to handle different shapes and more complex folding using a pair of pants. We designed two different scenarios: 1) folding one leg in half, where the overall shape differs from a rectangle, and 2) folding in half the pants with both legs already folded, where the shape is still rectangular but the cloth is already folded in multiple layers. The results, detailed in Table IV under the columns “Leg” and “Pants”, show AdaFold outperforming the Triangular baseline, confirming its versatility with various shapes and pre-folded configurations.
The overall outcome confirms the adaptability of our method, underscoring the benefit of using particle-based representations along feedback-loop manipulation to adapt to different object variations.
Figure 7: Visualization of: (top) the mask with no semantic tracked over time, (bottom) AdaFold upper mask tracked over time.
@@
Camera(s) No Semantic $\downarrow$ VLM $\downarrow$ AdaFold $\downarrow$
Front $0.39\pm 0.10$ $0.39\pm 0.11$ $0.09\pm 0.02$
Back $0.15\pm 0.07$ $0.15\pm 0.04$ $0.09\pm 0.05$
Both $0.21\pm 0.09$ $0.23\pm 0.08$ $\mathbf{0.04\pm 0.05}$
@@ VII-E Semantic Cloth Representation
Finally, we assess whether our proposed perception module improves the representation of the cloth. Unlike previous sections, the IoU is used as a quantitative description of the cloth state instead of a performance metric. As an evaluation dataset, we used a subset of the real-world trajectories recorded for the previous evaluations. We manually annotated the masks belonging to the upper or bottom half of the cloth to extract $P^{U}$ and $P^{B}$ to compute the ground truth IoU. To compare different state representations, we used the mean absolute error (MAE) computed between the estimated IoU from a specific cloth representation and the ground truth. Specifically, we compared the representations with semantic descriptors obtained through our perception module (AdaFold), against a baseline representation lacking semantic descriptors (No Semantic), and a segmentation method derived from a greedy selection of the highest confidence mask from the same VLM used for object segmentation at $t=0$ with prompts “upper half cloth”, “bottom half cloth” (VLM). We further integrated the comparison between one and two camera points of view. As shown in Table V, our approach always achieved the lowest MAE, with the best performance provided by leveraging the observations from both cameras. The representation with no semantic information obtained the worst results, confirming the challenges of distinguishing different folded states only relying on the observed points of the cloth. Integrating semantic information from the VLM did not yield good results, as the model always suggested the full cloth as a mask instead of the desired cloth regions. Trying to tune the prompts did not provide any improvement. In contrast, our method successfully segmented the desired cloth region. Fig. 7 shows qualitative results of the obtained masks. These results confirm that augmenting point clouds with semantic descriptors provides a better representation of different folded configurations.
@@ VIII Conclusion and Future Work
AdaFold framework leverages model-based feedback-loop manipulation to optimize cloth folding trajectories. It integrates semantic descriptors extracted from geometric features into point cloud representations. The experiments validated the hypothesis that AdaFold adapts folding trajectories to variations in physical properties, positions, and sizes of cloths. It was further showcased the potential of coupling a strong perception module with data-driven optimization strategies to perform feedback-loop manipulation. We plan to extend these results to different goal and reward configurations and a broader spectrum of manipulation tasks. We will investigate the integration of model-based and model-free learning approaches to address the computational demands of model-based planning.